The farm keeps 1000 Jersey cows and each produces 20 litres of milk that is sold at R5.00 per litre. About 100 litres of milk are consumed at the farmer's home while 1000 litres were donated to the orphanage. About 30 cows were sold at R5000 each while 15 cows were donated to the farm at R4950 for each cow. About 200 litres were stolen by two notorious farm workers.calculate gross income


1
Expert's answer
2021-05-12T07:26:30-0400
"1000\times20=20000"

"20000-100-1000-200=18700"

is total ptoduse milk.


"18700\times5=93500"

total rewenew from milk


"30\times5000+15\times4950=224250"

total rwenew from cows



"224250+93500=317750"